If a pyramid has a square base that measures 6 m on each side and a height of 5 m, then its volume is
Answer by jim_thompson5910(35256)

Volume of Square Pyramid = (1/3)*(Height)*(Base Length)^2


Volume of Square Pyramid = (1/3)*(5)*(6)^2


Volume of Square Pyramid = (1/3)*(5)*(36)


Volume of Square Pyramid = (1/3)*180


Volume of Square Pyramid = 60 cubic meters